# Orthogonal Blocking Kaczmarz Algorithm Based on Preprocessing Technology

### v2

@@ 5 Numerical experiments
In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK (k), GREBK (k), and POBK methods, verifying the effectiveness and efficiency of the algorithm. The matrices used for numerical experimental testing are all selected from 20 sparse matrices with application backgrounds in the $SuiteSparse\;Matrix\;Collection$, and have different scales (most of which are large sparse matrices). Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the system of equations. We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector.
@@
The numerical experimental results in Table 2 show that our proposed Kaczmarz (POBK) algorithm, which has been preprocessed and reordered into blocks, is much faster than the Kaczmarz (GRBK) and k-means direct block Kaczmarz (GRBK, GREBK (K), RBK (k)) algorithms of the random block algorithm in terms of iteration steps and computation time, and the improvement rate of algorithm iteration times and computation time is mostly above 90%. It is worth noting that for matrices in the complex field, such as the Kim1 matrix, the GRBK, GREBK (K), and RBK (k) algorithms encountered problems during partitioning and were unable to continue running. However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
@@
$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $POBK$ $Increase~{}Speed\left(\%\right)$
$Poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 $75$ $\left[99.81,99.90\right]$
CPU $40.63$ $43.43$ $28.71$ $0.15$ $\left[99.48,99.65\right]$
$Muu$ IT $606$ $176$ $269$ $1$ $\left[99.43,99.83\right]$
CPU $2.91$ $7.84$ $2.44$ $0.26$ $\left[89.34,96.68\right]$
$Poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 1 03410341\,0341 034 $\left[98.21,98.61\right]$
CPU $136.48$ $338.26$ $125.88$ $8.92$ $\left[92.91,97.36\right]$
$Ex29$ IT $253$ $10$ $116$ $8$ $\left[20.00,96.84\right]$
CPU $0.19$ $1.37$ $0.20$ $0.12$ $\left[58.33,91.24\right]$
$Kim1$ IT $NAN$ $NAN$ $NAN$ $52$ $\left[NAN,NAN\right]$
CPU $NAN$ $NAN$ $NAN$ $18.27$ $\left[NAN,NAN\right]$
$Bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 $296$ $\left[99.75,99.94\right]$
CPU $NAN$ $NAN$ $807.00$ $30.64$ $\left[96.20,98.42\right]$
$Qpband$ IT $627$ $335$ $292$ $1$ $\left[99.66,99.84\right]$
CPU $1.28$ $194.37$ $4.33$ $1.01$ $\left[21.09,99.48\right]$
$Rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 1 42214221\,4221 422 $\left[99.41,99.72\right]$
CPU $NAN$ $3094.14$ $3074.96$ $78.34$ $\left[95.17,97.47\right]$
$Obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $132$ $\left[97.37,99.65\right]$
CPU $158.43$ $1108.05$ $62.96$ $19.08$ $\left[69.70,98.28\right]$
$Spmsrtls$ IT $Inf$ $Inf$ $Inf$ $2$ $\left[99.99,99.99\right]$
CPU $NAN$ $NAN$ $NAN$ $2.05$ $\left[99.91,99.98\right]$
$Blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 1 38113811\,3811 381 $\left[93.78,99.62\right]$
CPU $456.49$ $32.55$ $261.30$ $10.66$ $\left[67.25,97.66\right]$
$Linverse$ IT $Inf$ $Inf$ $Inf$ $10$ $\left[99.99,99.99\right]$
CPU $NAN$ $NAN$ $NAN$ $0.54$ $\left[99.97,99.98\right]$
$Torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $132$ $\left[88.94,99.65\right]$
CPU $166.10$ $966.06$ $147.52$ $18.54$ $\left[87.43,98.08\right]$
$Polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 763363763363\,763363 763 5 12851285\,1285 128 $\left[98.59,98.97\right]$
CPU $NAN$ $1002.59$ $813.27$ $23.91$ $\left[97.06,98.05\right]$
$Jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 3 60336033\,6033 603 $\left[98.58,99.28\right]$
CPU $NAN$ $272.82$ $533.79$ $16.24$ $\left[94.05,97.78\right]$
$Dixmaanl$ IT 38 8753887538\,87538 875 $980$ 2 55825582\,5582 558 $9$ $\left[99.08,99.98\right]$
CPU $237.74$ $1687.74$ $17.05$ $0.63$ $\left[96.30,99.96\right]$
$Bcsstm39$ IT $20$ $20$ $20$ $1$ $\left[95.00,95.00\right]$
CPU $0.10$ $5189.52$ $19.14$ $0.04$ $\left[60.00,99.99\right]$
$Crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $1$ $\left[99.86,99.98\right]$
CPU $41.89$ $645.50$ $24.83$ $2.24$ $\left[90.98,99.65\right]$
$Chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $9$ $\left[99.75,99.90\right]$
CPU $3.88$ $5.08$ $2.03$ $0.05$ $\left[97.54,99.02\right]$
$Wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $160$ $\left[94.94,96.77\right]$
C PU $20.78$ $2193.32$ $36.76$ $20.57$ $\left[1.01,99.06\right]$
@@
Figure 5:  The number of iteration steps (left) and average computation time (right) for testing matrix $Poli3$ to reach convergence under different iteration methods.
Figure 6:  The number of iteration steps (left) and average computation time (right) for testing matrix $Polilarge$ to reach convergence under different iteration methods.
@@
Figure 7:  The number of iteration steps (left) and average computation time (right) for testing matrix $Jagmesh4$ to reach convergence under different iteration methods.
Figure 8:  The number of iteration steps (left) and average computation time (right) for testing matrix $Rajat07$ to reach convergence under different iteration methods.
@@
Figure 9:  The number of iteration steps (left) and average computation time (right) for testing matrix $Muu$ to reach convergence under different iteration methods.
5.1 The Number of matrix blocks
@@
When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm. The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$. Moreover, due to the differences in the results of each clustering of the k-means algorithm, the Kaczmarz algorithm, which is reflected in numerical experiments as a k-means clustering block, has significant differences in iteration time even for the same k-value, and the stability of the algorithm is not very good. In our proposed method, the selection of the number of blocks $k$ is also involved. Although it also needs to be set in advance, the value of $k$ can be determined based on the number of conditions in the matrix and the preprocessed ”width”. Numerical experiments have shown that the algorithm has strong stability and will not have significant differences in iteration time for the same $k$ value. Taking the $ex29$ matrix as an example, Figure 10 shows the iteration time changes of GRBK, RBK(k), GREBK(k), and POBK algorithms after 10 iterations with $k$ constant. It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in computation time, but also performs better in computation time stability.
@@
5.2 The Robustness of POBK Algorithm
Another advantage of our proposed algorithm is its robustness. Compared to GRBK, RBK(k), and GREBK(k) algorithms, even if different $k$ values are taken, the convergence rate of POBK algorithm does not change much, and it is still better than GRBK, RBK(k), and GREBK(k) algorithms. Taking $Poli$, $Muu$, $Crystm03$, $Torsion1$ matrices as examples, we compared the iteration time and steps required for convergence of GRBK, RBK(k), GREBK(k) algorithms with POBK algorithms at different values. The $k$ in POBK(k) represents the number of blocks, as detailed in Table 3 and Figure 11.
Table 3:   The IT and CPU required for the convergence of GRBK, RBK (k), GREBK (k) algorithms and POBK algorithms with different $k$ values.
$Poli$ $Muu$ $Crystm03$ $Torsion1$
$GRBK$ IT 78 9727897278\,97278 972 $716$ 4 $34$\,3434 343 37 4243742437\,42437 424
CPU $43.04$ $3.38$ $44.20$ $166.10$
$RBK(k)$ IT 36 2233622336\,22336 223 $213$ $863$ 1 19411941\,1941 194
CPU $27.10$ $11.56$ $860.76$ $966.06$
$GREBK(k)$ IT 51 6835168351\,68351 683 $410$ 1 03010301\,0301 030 5 04350435\,0435 043
CPU $27.60$ $3.32$ $16.48$ $147.52$
$POBK(3)$ IT $75$ $1$ $1$ $108$
CPU $0.16$ $0.24$ $2.30$ $34.07$
$POBK(5)$ IT $165$ $7$ $9$ $114$
CPU $0.36$ $0.62$ $7.56$ $26.76$
$POBK(10)$ IT $250$ $8$ $11$ $9122$
CPU $0.80$ $0.55$ $8.29$ $24.21$
$POBK(15)$ IT $792$ $8$ $16$ $132$
CPU $3.16$ $0.49$ $9.82$ $18.54$
$POBK(20)$ IT $851$ $9$ $10$ $151$
CPU $4.27$ $0.52$ $5.75$ $20.47$
Figure 11:  Comparison chart of $CPU$ required for convergence of $GRBK$, $RBK(k)$, $GREBK(k)$ algorithms and $POBK$ algorithms with different $k$ values.
@@ 6 Conclusions
This article proposes a new preprocessed orthogonal block wise Kaczmarz (POBK) algorithm based on random block Kaczmarz and k-means technology. We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate block diagonal matrix. Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and calculate the inner product between the centroid coordinates to measure the orthogonality between different blocks, so that we can group the orthogonal blocks into one class. We conducted theoretical analysis on the proposed algorithm, proved the convergence of the POBK algorithm, and provided a theoretical analysis of its faster convergence. The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), and GREBK(k) algorithms in terms of iteration steps and CPU time. The preprocessing method makes the original matrix structure easier to handle, and the new partitioning method allows for two orthogonal projections on two mutually orthogonal hyperplanes during each iteration, which makes the iterative solution quickly converge to the exact solution. At present, existing preprocessing methods include normalizing and orthogonalizing matrices, using preconditioners to preprocess matrices, and the partitioning idea mainly uses k-means clustering for partitioning. We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition matrices, in order to improve the numerical performance of the POBK algorithm. This is a valuable work worth studying in the future.

### v3

@@ 5 Numerical experiments
5.1 Numerical simulation
In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK(k), GREBK(k), aRBK and POBK methods, verifying the effectiveness and efficiency of the algorithm. The matrices used for numerical experimental tests are 20 sparse matrices selected from the $SuiteSparse\;Matrix\;$ $Collection$ with application background and varying in size, condition number and density. Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the linear system. We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector space.
@@ 5.1 Numerical simulation
Since the GRBK, GREBK(k) and RBK(k) algorithms involve only one block per iteration, yet the aRBK and POBK algorithms involve all the blocks of the matrix per iteration, a direct comparison of the number of iterative steps (IT) is not fair. As can be seen from Table 3, even with $k$ times the number of iteration steps of the POBK algorithm($k<20$), it is still far superior to the GRBK, GREBK(k) and RBK(k) algorithms. As for the computation time, the POBK algorithm is much faster than the randomized chunking algorithms Kaczmarz (GRBK, aRBK) and the k-means direct chunking Kaczmarz (GREBK(k), RBK(k)) algorithms, and we give a comparison of the average running time of the different algorithms for some of the matrices in Figures 5, 6 and 7. It is worth noting that for the matrices over complex field, such as kim1 matrices, the GRBK, GREBK(k), and RBK(k) algorithms encountered problems during the chunking process and could not continue to run. However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
From the convergence analysis in Section 4, it can be seen that the convergence speed of the POBK algorithm is related to the maximum and minimum non-zero singular values of the blocks and the angle of the hyperplane. We analyzed the matrix structure, and found that other matrices either have smaller ”bandwidth” after preprocessing to accelerate the convergence of the algorithm, or the gap between the largest and smallest non-zero singular values after preprocessing the chunks becomes smaller to accelerate the convergence of the algorithm. However, wathen100 is a very special matrix. Its original matrix is presented as a banded form, and the ”bandwidth” of the RCM algorithm after preprocessing is slightly larger than the original matrix bandwidth (see Figure 8). At the same time, the difference between the maximum and the minimum non-zero singular values after the chunking is also larger (the minimum is 6, and the maximum can reach 40), which will undoubtedly also make the POBK algorithm converge slowly. As a result, the wathen100 matrix has a smaller number of iterations, but the iteration time is comparable to that of the GRBK algorithm. In addition, we find that there is almost no difference in the running time of the wathen100 matrix between preprocessing with the RCM algorithm and without reordering, which reinforces our idea. In this case (sparse matrices with a narrow banding structure), using the original matrix directly may be a better choice. This is also in line with our theme that the POBK algorithm presents good convergence for sparse matrices with scattered arrangements.
@@
$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $POBK$
$poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 1 72117211\,7211 721 $28$
CPU $40.63$ $43.43$ $28.71$ $33.75$ $0.07$
$muu$ IT $606$ $176$ $269$ $13$ $1$
CPU $2.91$ $7.84$ $2.44$ $3.51$ $0.22$
$poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 $899$ 1 03410341\,0341 034
CPU $136.48$ $338.26$ $125.88$ $121.63$ $7.38$
$ex29$ IT $253$ $10$ $116$ $7$ $8$
CPU $0.19$ $1.37$ $0.20$ $0.51$ $0.10$
$kim1$ IT $NAN$ $NAN$ $NAN$ $124$ $52$
CPU $NAN$ $NAN$ $NAN$ $299.93$ $18.27$
$bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 18 3251832518\,32518 325 $112$
CPU $NAN$ $NAN$ $807.00$ $3574.54$ $12.97$
$qpband$ IT $627$ $335$ $292$ $11$ $1$
CPU $1.28$ $194.37$ $4.33$ $3.36$ $1.01$
$rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 6 89668966\,8966 896 1 08010801\,0801 080
CPU $NAN$ $3094.14$ $3074.96$ $944.40$ $57.52$
$tols4000$ IT $491$ $10$ $927$ $11$ $1$
CPU $0.29$ $4.63$ $0.49$ $0.25$ $0.04$
$obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $339$ $127$
CPU $158.43$ $1108.05$ $62.96$ $195.34$ $18.72$
$spmsrtls$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
CPU $NAN$ $NAN$ $NAN$ $NAN$ $2.24$
$blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 2 30823082\,3082 308 1 05510551\,0551 055
CPU $456.49$ $32.55$ $261.30$ $151.08$ $10.10$
$linverse$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
CPU $NAN$ $NAN$ $NAN$ $NAN$ $0.42$
$torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $386$ $132$
CPU $166.10$ $966.06$ $147.52$ $209.02$ $18.57$
$polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 $63 7$ 69776977\,6977 697 5 12851285\,1285 128
CPU $NAN$ $1002.59$ $813.27$ $975.98$ $20.40$
$jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 7 10471047\,1047 104 2 87328732\,8732 873
CPU $NAN$ $272.82$ $533.79$ $292.11$ $10.40$
$bcsstm39$ IT $20$ $20$ $20$ $4$ $1$
CPU $0.10$ $5189.52$ $19.14$ $4.09$ $0.04$
$crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $29$ $1$
CPU $41.89$ $645.50$ $24.83$ $29.81$ $2.15$
$chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $20$ $9$
CPU $3.88$ $5.08$ $2.03$ $0.61$ $0.04$
$wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $49$ $150$
C PU $20.78$ $2193.32$ $36.76$ $30.46$ $20.57$
@@
Figure 5:  Average computational time for test matrices $poli3$ and $polilarge$ to reach convergence for different iteration methods.
@@
Figure 6:  Average computational time for test matrices $jagmesh4$ and $rajat07$ to reach convergence for different iteration methods.
@@
Figure 7:  Average computational time for test matrices $muu$ and $chem97ztz$ to reach convergence for different iteration methods.
Figure 8:   Comparison of matrix wathen100 before and after RCM preprocessing.
5.2 The Number of matrix blocks $k$
When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm. According to Equation (4.14), the convergence speed of the POBK algorithm is directly related to the cosine of the angle between row blocks. When the number of dimensions $m$ increases, $m>2$. Here, the pinch angle $\theta$ involves both mutually orthogonal blocks and non-mutually orthogonal blocks. When we place mutually orthogonal blocks in a category, the cosine of the angle between these mutually orthogonal blocks will be very small or even close to 0. At this point, it is important to pay attention to the non-mutually orthogonal blocks, i.e., the number of nonzero elements in the cosine table, and the size of these nonzero elements for the rate of convergence of the POBK algorithm to be critical, which also determines the number of matrix chunks. Here we define two variables: the Orthogonality Proportion $zn$ and the Non-orthogonality Severity $nn$ to serve as selection criteria for the number of chunks.
Definition 1. For the cosine value table $C$, if the total number of elements is $num=k\times k$, the number of zero elements is $n_{1}$, then the Orthogonality Proportion $zn$ is denoted by
$zn=\frac{n_{1}}{num}=\frac{n_{1}}{k\times k}.$
Definition 2. For the cosine value table $C$, the total number of elements is $num=k\times k$, the number of non-zero elements is $n_{2}$, and the average of these non-zero elements is $nm$, then the Non-orthogonality Severity $nn$ is denoted by
$nn=\frac{n_{2}\times nm}{num}=\frac{n_{2}\times nm}{k\times k}.$
It is worth noting that the variable $nn$ is not directly defined by the number of non-zero elements in the matrix $C$ as a percentage of the total number of elements, because we have to consider not only the number of non-zero elements, but also the size of the values of these non-zero elements. If these non-zero elements have large values, it means that the angle between some blocks is small, which also makes the iteration slower, which is exactly what we do not want to see.
We take the ex29 and wathen100 matrices as examples. For the ex29 matrix, from Figure 9, $zn$ increases as $k$ increases, and $nn$ is small at $k=2,3,5$, but the value of $zn$ at $k=2,3$ is too small, so it is appropriate to choose $k=5$, and the lowest point of the running time box-and-line diagram is also at $k=5$ from Figure 11. For the wathen100 matrix, from Figure 10, $zn$ obtains its maximum value at $k=2$, and it happens that $nn$ also obtains its minimum value at $k=2$, so the choice of $k=2$ is appropriate, and the lowest point of the running time box-and-line plot is also at $k=2$ from Figure 11.
Figure 9: $zn$ and $nn$ values of the ex29 matrix for different values of $k$.
Figure 10: $zn$ and $nn$ values of the muu matrix for different values of $k$.
Figure 11:  Boxplots of runtime for ex29 and muu matrices for different values of $k$.
5.3 Stability of the POBK algorithm
@@
The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$. The difference in the results of each clustering of the k-means algorithm is reflected in numerical experiments as the block Kaczmarz algorithm for k-means clustering, where there is a significant difference in the iteration time even with the same value of $k$, which can lead to the algorithm’s stability is not very good. In our proposed method, given a test matrix $A$, which is preprocessed by the RCM algorithm as an approximate banded matrix $\widetilde{A}$, and $\widetilde{A}$ is unique, and since we perform sequential uniform chunking, the form of the chunks is fixed once $k$ is determined, and thus the algorithm is stable for the same $k$ value. Numerical experiments also show that the algorithm is stable and there is no significant difference in the iteration time for the same value of $k$. Taking the ex29 matrix as an example, Figure 12 shows the iteration time changes of GRBK, RBK(k), GREBK(k), aRBK and POBK algorithms after 10 iterations with the same $k$. It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in terms of computation time, but also performs better in terms of stability. The stronger stability of the POBK algorithm can also be seen from the POBK runtime boxplot in Section 5.3.
@@ 6 Conclusions
In this paper, we propose a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques. We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate banded matrix. Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and the cosine of the angle between the center of mass coordinates is computed to measure the orthogonality between the different blocks, so that the orthogonal blocks can be grouped together. We conducted theoretical analysis on the proposed algorithm, proved the convergence of the POBK algorithm, and provided a theoretical analysis of its faster convergence. The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), GREBK(k), and aRBK algorithms in terms of CPU time. The preprocessing method makes the original matrix structure easier to handle, and the new partitioning method allows for two orthogonal projections on two mutually orthogonal hyperplanes during each iteration, which makes the iterative solution quickly converge to the exact solution. At present, existing preprocessing methods include normalizing and orthogonalizing matrices, using preconditioners to preprocess matrices, and the partitioning idea mainly uses k-means clustering for partitioning. We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition coefficient matrices, in order to improve the numerical performance of the POBK algorithm. This is a valuable work worth studying in the future.
@@
CONFLICT OF INTEREST STATEMENT
This study does not have any conflicts to disclose.
DATA AVAILABILITY STATEMENT
Our computational data comes from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). These data can be freely downloaded and analyzed.